### v2

# Tilings from Tops of Overlapping Iterated Function Systems

Abstract.
The top of the attractor of a hyperbolic iterated function system $\left\{f_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}|i=1,2,\dots,M\right\}$ is defined and used to extend self-similar tilings to overlapping systems. The theory provides sequences of approximate supertiles that converge to tilings. Individual tiles in a tiling are limits of nested decreasing sequences of approximate tiles. Examples include systems of finite type, tilings related to aperiodic monotiles, and ones where there are infinitely many distinct but related prototiles.
1. Introduction
How can self-similar tiling theory [3, 15] be extended to collections of overlapping tiles?
Tilings associated with iterated function systems (IFS) that obey the open set condition (OSC) or are of finite type have been investigated in other papers, see for example [4, 5, 7, 9] and references therein.
Here we use top addresses, defined in Section 2.2, to extend self-similar tiling theory to include tilings generated by iterated function systems (IFS) that do not obey the open set condition and are not of finite type. The tiles in these tilings are closures of the limits of decreasing sequences of approximate tiles and have nonempty interiors. Individual tiles become fully defined after finitely many steps, while the tilings themselves become fully defined only in the infinite limit. Since our construction depends mainly on topology rather than geometry, it also applies to point-fibred IFSs acting on compact Hausdorff spaces, see [7, 16].
In [6] top addresses were introduced to define tilings in special cases. This paper elucidates the situation in [6] and answers in the affirmative the conjecture at the end of that paper. A distinct approach to overlapping tilings has been provided by Akiyama and co-workers [2].
As an application, we show how our construction applies in the context of aperiodic monotiles [19].
2. Iterated Function Systems, Tops, Blowups, and Tilings
2.1. Iterated function systems
Let $F$ be an iterated function system (IFS)
[…]
$d(f_{i}(x),f_{i}(y))\leq\lambda d(x,y),\text{ some }0<\lambda<1,\text{ for all }x,y,i.$
Then it is known [14] that $F$ possesses a unique attractor, the only non-empty compact subset of $\mathbb{R}^{n}$ which obeys
$A=\bigcup\limits_{i=1}^{M}f_{i}(A).$
We will use the fact that if $A$ has non-empty interior, then it is the closure of its interior, $A=\overline{A^{\circ}}.$
We will need the following notions related to symbolic handling of subsets of $A.$ Let $\mathbb{N}$ be the positive integers. Consider $\left\{1,2,\dots,M\right\}^{\mathbb{N}}$, the set of infinite strings of the form $\mathbf{j}=j_{1}j_{2}\dots$ where each $j_{i}$ belongs to $\left\{1,2,\dots,M\right\}$. We define $i:\left\{1,2,\dots,M\right\}^{\mathbb{N}}\rightarrow\left\{1,2,\dots,M\right\}^{\mathbb{N}}$ for $i\in\{1,2,\dots,M\}$ by $i(\mathbf{j)=}ij_{1}j_{2}\dots$ We may also write $k_{1}k_{2}\dots k_{l}\mathbf{j}$ to mean the string $k_{1}k_{2}\dots k_{l}j_{1}j_{2}\dots.$ The string $\mathbf{j}$ truncated to length $m$ is denoted by $\mathbf{j}|m=j_{1}j_{2}\dots j_{m},$ and we define
$f_{\mathbf{j}|m}=f_{j_{1}}f_{j_{2}}\dots f_{j_{m}}=f_{j_{1}}\circ f_{j_{2}}\circ\dots\circ f_{j_{m}}.$
We define a metric $d^{\prime}$ by $d^{\prime}(\mathbf{j},\mathbf{k})=2^{-\max\{n|j_{m}=k_{m},m=1,2,...,n\}}$ for $\mathbf{j}\neq\mathbf{k}$, so that $\left(\left\{1,2,\dots,M\right\}^{\mathbb{N}},d^{\prime}\right)$ is a compact metric space.
Then a continuous surjection $\pi:\left\{1,2,\dots,M\right\}^{\mathbb{N}}\rightarrow A$ is defined by
$\pi(\mathbf{j})=\lim_{m\rightarrow\infty}f_{\mathbf{j}|m}(x)=\lim_{m\rightarrow\infty}f_{j_{1}}f_{j_{2}}\dots f_{j_{m}}(x).$
It is well-known [14] that the limit is independent of $x$. Also, the convergence is uniform in $\mathbf{j}$ over $\left\{1,2,\dots,M\right\}^{\mathbb{N}},$ and uniform in $x$ over any compact subset of $\mathbb{R}^{n}$. We say $\mathbf{j}\in\Sigma$ is an address of the point $\pi(\mathbf{j})\in A$. Each point in $A$ has at least one address.
2.2. Tops
Since $\pi:\left\{1,2,\dots,M\right\}^{\mathbb{N}}\rightarrow A$ is continuous and onto, it follows that $\pi^{-1}(x)$ is closed for all $x\in A.$ Hence, a map $\tau:A\rightarrow\left\{1,2,\dots,M\right\}^{\mathbb{N}}$ and a set $\Sigma$ are well-defined by
$\tau(x)$ $:=\max\{\mathbf{k\in}\Sigma|\pi(\mathbf{k)=}x\},$
$\Sigma$ $:=\tau(A)=\left\{\tau(x):x\in A\right\},$
where the maximum is with respect to lexicographical ordering. We call $\tau(x)$ the top address of $x\in A$.
The top address $\tau(x)$ of $x\in A$ and the set of top addresses $\Sigma$ may be calculated by following orbits under the dynamical system $D:A\rightarrow A$ as follows. Partition $A$ into $\left\{A_{i}\right\}$ according to
$A_{1}=f_{1}(A),A_{2}=f_{2}(A)\backslash A_{1},\ A_{3}=f_{3}(A)\backslash(A_{1}\cup A_{2}),\dots,A_{M}=f_{M}(A)\backslash\cup_{n\neq M}A_{n}.$
We always assume $A_{M}\neq\emptyset.$ Define the orbit $\left\{x_{n}\right\}_{n=1}^{\infty}$ of $x=x_{1}\in A,$ by $x_{n+1}=Dx_{n}=f_{i_{n}}^{-1}(x_{n})$ where $i_{n}$ is the unique index such that $x_{n}\in A_{i_{n}}$. Then $\tau(x)=i_{1}i_{2}i_{3}\dots\in\Sigma$.
Top addresses are related to $\beta$-expansions, see in particular [21]. See also [1] and references therein.
Define the critical set of $F$ to be
$C=\overline{{\textstyle\bigcup\limits_{i}}\partial A_{i}\backslash\partial A}.$
Let $\sigma:\{1,2,\dots,M\}^{\mathbb{N}}\rightarrow\{1,2,\dots,M\}^{\mathbb{N}}$ be the shift operator defined by $\sigma(\mathbf{j)=}j_{2}j_{3}\dots$. Then
$f_{i}\circ\pi=\pi\circ i\mathbf{\ }\text{and }f_{j_{1}}^{-1}\circ\pi\left(\mathbf{j}\right)=\pi\circ\sigma\left(\mathbf{j}\right)\text{ for all }i\text{ and }\mathbf{j}.$
Key to the results in this paper is the total shift invariance of $\Sigma$, as follows.
Lemma 1.
$\Sigma=\sigma\left(\Sigma\right)$.
Proof.
This is readily checked [6]. ∎
2.3. Tops at finite depth
Define $\Sigma_{n}$ to be the elements of $\Sigma$ truncated to length $n\in\mathbb{N}$,
$\Sigma_{n}=\left\{\left(\mathbf{j}|n\right)|\mathbf{j}\in\Sigma\right\}.$
Corresponding to $\mathbf{i}|n\in$ $\Sigma_{n}$, define members of a partition of $A$ at depth $n$ by
$A_{\mathbf{i}|n}=\left\{\pi(\mathbf{j}):\mathbf{j\in}\Sigma,\left(\mathbf{j|}n\right)=\left(\mathbf{i|}n\right)\right\}.$
Equivalently,
[…]
We will need the following lemmas.
Lemma 2.
Let $\mathbf{k}\in\Sigma.$ Then $f_{k_{1}}\left(A_{(\sigma\left(\mathbf{k}\right)\mathbf{)|}n-1}\right)\supseteq A_{\mathbf{k|}n}$ for all $n\in\mathbb{N}$, $n>1$.
Proof.
We compare the sets
$\{f_{k_{1}...k_{n}}(x)|f_{k_{1}...k_{n}}(x)\notin f_{l_{1}l_{2}...l_{n}}(A)\text{ for all }l_{1}...l_{n}>k_{1}...k_{n}\}$
and
$\{f_{k_{1}...k_{n}}(x)|f_{k_{1}}f_{k_{2}...k_{n}}(x)\notin f_{k_{1}}f_{l_{2}...l_{n}}(A)\text{ for all }l_{2}...l_{n}>k_{2}...k_{n}\}.$
The condition in the latter expression is less restrictive. ∎
Also, since $\Sigma$ is shift invariant, we have:
Lemma 3.
Let $n>1.$ If $i_{1}i_{2}\dots i_{n-1}i_{n}\in\Sigma_{n}$, then both $i_{2}\dots i_{n-1}i_{n}\$and $i_{1}i_{2}\dots i_{n-1}\$belong to $\Sigma_{n-1}.$
2.4. Top blowups and tilings
We focus on the overlapping case, where the OSC does not hold, with particular interest in the case where $A$ has nonempty interior. We show how tilings of $\mathbb{R}^{n}$ can be defined, generalizing well known constructions that apply when the OSC holds [4, 7, 20].
Let $\overleftarrow{\Sigma}\subset\{1,2,\dots,M\}^{\mathbb{N}}$ be the set strings $\mathbf{l}=l_{1}l_{2}\dots$ such that $l_{n}l_{n-1}\dots l_{1}\in\Sigma_{n}$ for all $n.$ Note that $\overleftarrow{\Sigma}$ is closed but $\Sigma$ may not be. We define, for $\mathbf{j\in}\overleftarrow{\Sigma}$ and $k\in\mathbb{N}$,
$f_{-\mathbf{j}|k}=f_{j_{1}}^{-1}f_{j_{2}}^{-1}\dots f_{j_{k}}^{-1}=f_{j_{1}}^{-1}\circ f_{j_{2}}^{-1}\circ\dots\circ f_{j_{k}}^{-1}.$
Strichartz [20] defines the blowup $\mathcal{A}\left(\mathbf{j}\right)$ of $A$ corresponding to $\mathbf{j\in}\overleftarrow{\Sigma}$ by
$\mathcal{A}\left(\mathbf{j|}n\right)=\bigcup\limits_{l=1}^{n}f_{-\mathbf{j}|l}\left(A\right)\text{ and }\mathcal{A}\left(\mathbf{j}\right)=\bigcup\limits_{n=1}^{\infty}\mathcal{A}\left(\mathbf{j|}n\right).$
(Note that Strichartz was concerned with situations where the OSC holds, in which case $\overleftarrow{\Sigma}=\{1,2,\dots\}^{\mathbb{N}}.$) Here the unions are of increasing nested sequences of sets so $\mathcal{A}\left(\mathbf{j|}n\right)=f_{\mathbf{j}|n}^{-1}\left(A\right).$ Note that $\mathcal{A}\left(\mathbf{j|}n\right)$ is related to $\mathcal{A}\left(\mathbf{i|}n\right)$ by the isometry $\left(f_{-\mathbf{j}|n}\right)\left(f_{-\mathbf{i}|n}\right)^{-1}$. But possible relationships between $\mathcal{A}\left(\mathbf{i}\right)$ and $\mathcal{A}\left(\mathbf{j}\right)$ are subtle because inverse limits are involved.
Under conditions on $F$ and fixed $\mathbf{j}\in\overleftarrow{\Sigma}$, stated in Theorem 1, we construct a collection of sets as follows. We define:
(2.1) $\Pi(\mathbf{j}|k)$ $:=\left\{f_{-\mathbf{j}|k}(A_{\mathbf{t|}k+1})|\mathbf{t\in}\Sigma\right\},$
[…]
For $\mathbf{t\in}\Sigma$ and $\mathbf{j}\in\overleftarrow{\Sigma},$ we may refer to sets of the form $f_{-\mathbf{j}|k}(A_{\mathbf{t|}k+1})$ and $f_{-\mathbf{j}|k}(\overline{A_{\mathbf{t|}k+1}}),$ where the bar denotes closure, as pre-tiles and tiles respectively.
Figure 1. The right hand panel represents the attractor of the IFS in Equation 3.1 partitioned at depth one into four pre-tiles, indicated by different colours. Dotted lines indicate open boundaries. The left hand panel is similar, but partitioned to depth two.
Figure 2. The maps $f_{4}^{-1}$ (left) and $f_{1}^{-1}$ (right) have been applied to the right-hand partition in Figure 1. Notice the different positionings of the closed and open edges of the pre-tiles that cover $\left(0,1\right)\times\left(0,1\right).$
When the OSC holds and $A$ has nonempty interior, that is $A^{\circ}\neq\emptyset$, the limit in Equation 2.1 is well defined because $\overline{A_{\mathbf{t|}k+1}}=f_{\mathbf{t|}k+1}(A)$ so that, up to closure of the sets $A_{\mathbf{t|}k+1}$, we have $\Pi(\mathbf{j}|k)\subset\Pi(\mathbf{j}|k+1)$ for all $k,$ so $(\left\{f_{-\mathbf{j}|k}(\overline{A_{\mathbf{t|}k+1}})|\mathbf{t\in}\Sigma\right\})_{k=1}^{\infty}$ is an increasing nested sequence of finite collections of just-touching sets. In this case, at each level new pre-tiles are added, while the tiles defined by taking the closures of pre-tiles belonging to the previous levels remain unchanged. In Section 3.1 we provide a simple example of pre-tiles and use of the notion of the top of an attractor. Tilings constructed in this manner, though with different language, are discussed in [8, 10].
But the general situation is quite different because nested sequences of pre-tiles may change substantially from one level to the next. As we will show, the collection of sets $\Pi(\mathbf{j})$ is well defined in situations where $\Pi(\mathbf{j}|k)$ may not be a subset of $\Pi(\mathbf{j}|k+1)$, even when the closures of the pre-tiles is taken. In general, for $\mathbf{j}\in\overleftarrow{\Sigma}$, here is an informal preview of what happens. At each successive value of $k$, the pre-tiles from the previous level are smaller, and new pre-tiles are added. Sequences of sucessively smaller pre-tiles cease to change after finitely many steps, and their closures remain thereafter as established tiles with nonempty interiors, members of a tiling that is fully defined only in the limit as $k\rightarrow\infty$. Two simple examples of this process, illustrating our notion of pre-tiles and showing how pre-tiles may be successively smaller but finally stabilize, are given in Section 3. The reader may prefer to consult Section 3 before proceeding to our main result, Theorem 1.
In the overlapping case it is not obvious that the pre-tiles we will define have non-empty interiors. We will need the following Lemma.
Lemma 4.
[…]
Proof.
[…]
In Theorem 1 we consider expressions of the form
$S(l,\mathbf{i},k):=f_{j_{1}}^{-1}\dots f_{j_{k}}^{-1}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})$
where $k,l\geq 0$ and $\mathbf{i\in}\Omega_{l}$ with
$\Omega_{l}:=\left\{\mathbf{i}\in\Sigma:j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}\in\Sigma_{k+1}\text{ for all }k\geq l+1\right\}.$
Sets of the form $S(l,\mathbf{i},k)$ are supersets of pre-tiles that occur in the expression $\lim_{k\rightarrow\infty}\Pi(\mathbf{j}|k)$. The key observation underlying Theorem 1 is that, under a condition on $\mathbf{j}$, as $k$ increases the sequence of sets $S(l,\mathbf{i},k)$ decreases for finitely many steps after which it does not change.
Theorem 1.
Let $F$ be a strictly contractive IFS on $\mathbb{R}^{n}$ whose attractor $A$ has nonempty interior. Let $C$ be the critical set of $F.$ Let $\mathbf{j\in}\overleftarrow{\Sigma}$, $\widehat{\lambda}\in(\lambda,1),$ $x_{0}\in A,$ and $\varepsilon>0,$ be fixed such that the sequence $\left\{x_{n}=f_{j_{n}}f_{j_{n-1}}\dots f_{j_{1}}(x_{0})\right\}_{n=1}^{\infty}$ obeys $d(x_{n},C)>$ $\varepsilon\left(\widehat{\lambda}\right)^{n}$ for all $n$.
(i) For fixed $l\in\mathbb{N}$ and $\mathbf{i\in}\Omega_{l}$, the sequence $\left\{S(l,\mathbf{i},k):k=0,1,\dots\right\}$ is nested decreasing according to $S(l,\mathbf{i},0)\supseteq S(l,\mathbf{i},1)\supseteq\dots$ and there is finite $K$ such that the sequence converges in a finite number of steps to a set $S(l,\mathbf{i})$ with non-empty interior
$S(l,\mathbf{i}):=S(l,\mathbf{i},K)=S(l,\mathbf{i},K+1)=\dots.$
(ii) Let
$S(l):=\{S(l,\mathbf{i}):\mathbf{i}\in\Omega_{l}\}.$
Then $\left\{S(l)\right\}_{l=0}^{\infty}$ is a nested increasing sequence of collections of sets with non-empty interiors, according to
$S(0)\subset S(1)\subset\dots.$
(iii) Define $\Pi(\mathbf{j})$ as the union limit of the above sequence, that is
$\Pi(\mathbf{j}):={\textstyle\bigcup\limits_{n=1}^{\infty}}S(n).$
Then $\Pi(\mathbf{j})$ provides a tiling of a region that includes
$f_{j_{1}}^{-1}\circ f_{j_{2}}^{-1}\circ\dots\circ f_{j_{l}}^{-1}\circ f_{j_{l+1}}^{-1}(A_{j_{l+1}})$
for all sufficiently large $l$. In particular, if $\mathbf{j}$ includes the symbol $1$ infinitely often, then $\Pi(\mathbf{j})$ provides a tiling of a region that includes the blowup $\mathcal{A}\left(\mathbf{j}\right)$.
Proof.
[…]
To prove the first part of (i), observe that by Lemma 2,
$f_{j_{k+1}}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})\supseteq A_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}}$
and apply $f_{j_{1}}^{-1}\dots f_{j_{k+1}}^{-1}$ to both sides.
For the remainder of (i), notice that if $f_{j_{k+1}}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})\cap C=\emptyset,$ then for any $f_{\omega_{1}\omega_{2}\dots\omega_{k+1}}(A)$ intersecting $f_{j_{k+1}}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})$ such that $\omega_{1}\omega_{2}\dots\omega_{k+1}>j_{k+1}j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}$, we must have $\omega_{1}=j_{k+1}$. That is, $f_{j_{k+1}}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}})=A_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}},$ and by applying $f_{j_{1}}^{-1}f_{j_{2}}^{-1}\dots f_{j_{k+1}}^{-1}$ to both sides we get $S(l,\mathbf{i,}k)=S(l,\mathbf{i,}k+1).$ So it suffices to find a large enough $K$ such that $A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}}$ contains no points of $C$ for all $k\geq K$.
But $A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}}$ is contained in $f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}}(A)$ and for any fixed $l$ and $k$ sufficiently larger than $l,$ the condition on $\mathbf{j\in}\overleftarrow{\Sigma}$, $\widehat{\lambda}\in(\lambda,1),$ $x_{0}\in A,$ and $\varepsilon>0,$ ensures that $f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}}(A)$ does not meet $C.$ Indeed,
(2.2) $d(f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}l+1\dots j_{1}}(x_{0}),f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}}(A))\leq\lambda^{k-l}|A|,$
where $\left|A\right|=\max\left\{d(x,y)|x,y\in A\right\}$ with $d$ the metric on $\mathbb{R}^{n},$ and
(2.3) $d(f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}\dots j_{1}}(x_{0}),C)\geq\left(\widehat{\lambda}\right)^{k+1}\varepsilon.$
So the shortest distance between $f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}}(A)$ and $C$ is greater than
(2.4) $\left(\widehat{\lambda}\right)^{k+1}\varepsilon-\lambda^{k-l}|A|$
which is strictly positive, for any fixed $l$ and any $k$ that is sufficienly larger than $l.$
To prove (ii), namely that $S(l)\subset S(l+1),$ note that
$S(l)$ $=\{S(l,\mathbf{i}):\mathbf{i}\in\Omega_{l}\}$
$=\{f_{j_{1}}^{-1}\dots f_{j_{k}}^{-1}(A_{j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}}):\text{for all }k\text{ as in (i)},\text{ }\mathbf{i}\in\Omega_{l}\},$
while
$S(l+1)$ $=\{S(l+1,\mathbf{i}):\mathbf{i}\in\Omega_{l+1}\}$
$=\{f_{j_{1}}^{-1}\dots f_{j_{k}}^{-1}(A_{j_{k}\dots j_{l+3}j_{l+2}i_{1}\dots i_{l+2}}):\text{for all }k\text{ as in (i)},\mathbf{i}\in\Omega_{l+1}\}.$
Fix $k$ sufficiently large, say $k\geq K,$ so that both expressions on the right hold. Then, using Lemma 3, the set of allowed indices in the first expression, ones with tails of the form $j_{l+1}i_{1}\dots i_{l+1},$ where $l+1$ is fixed, is a subset of the set of allowed indices in the second expression, with tails of the form $i_{1}\dots i_{l+2}$. So $S(l)\subset S(l+1)$. This completes the proof of (ii).
To prove (iii) we make the following observations. (a) Let $K_{l}$ be the maximum value of
$K_{l}:=\min\{k>l:f_{j_{k+1}j_{k}\dots j_{l+2}j_{l+1}i_{1}\dots i_{l+1}}(A)\cap C=\varnothing\text{ }\forall\text{ }\mathbf{i}\in\Omega_{l}\}.$
Then it follows from (i) that the $k>l+1$ in the definition of $\Omega_{l}$ can be replaced by $l+1\leq k\leq K_{l}.$ (b) It can be seen from Equation 2.4 that $\left\{K_{l}\right\}$ is a non-increasing. In fact, upon expansion, for
(2.5) $l>(\log_{\widehat{\lambda}/\lambda}(|A|/\varepsilon\lambda)-1)/(1+\log_{\widehat{\lambda}/\lambda}(\lambda))$
we have $K_{l}=l+1$.
It follows that
${\textstyle\bigcup}S(l)$ $={\textstyle\bigcup\limits_{i\in\Omega_{l}}^{\infty}}f_{j_{1}}^{-1}\dots f_{j_{K_{l}}}^{-1}(A_{j_{K_{l}}\dots j_{l+1}i_{1}\dots i_{l+1}})$
$=f_{-\mathbf{j}|l}(f_{-\sigma^{l}(\mathbf{j)|}K_{l}-l}(A_{j_{K_{l}}\dots j_{l+1}}))\text{ by (a)}$
[…]
The penultimate statement in the theorem follows on noting that $f_{1}^{-1}(A_{1})=A$ and that $\left\{f_{-\mathbf{j}|l}(A)\right\}_{l=1}^{\infty}$ is a nested increasing sequence. In particular, the sequence contains $A,$ and in particular the fixed-point of $f_{1},$ for all large enough $l,$ and is acted on by $f_{1}^{-1}$ infinitely many times, so it contains $f_{1}^{-1}(A),$ $f_{1}^{-1}\circ f_{1}^{-1}(A),\dots.$
Example 1.
The tiling $\Pi(\overline{1})$ is well-defined and its support is the blowup ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(A).$ In this case the set of tiles is exactly ${\textstyle\bigcup\limits_{k}}f_{1}^{-k}(\{f_{\mathbf{i}|(k+1)}(A):$ $\mathbf{i\in}\Sigma\mathbf{\},}$ see also [6].
Example 2.
If $\mathbf{j\in}\overleftarrow{\Sigma}$ is periodic and contains the symbol $1,$ then the tiling $\Pi(\sigma^{m}\mathbf{j})$ is well-defined for all $m=0,1,2\dots$.
Example 3.
If $\mathbf{j\in}\overleftarrow{\Sigma}$ contains the symbol $k$ infinitely often, and $A_{k}$ contains the fixed point of $f_{k}$ in its interior, then $\Pi(\mathbf{j})$ tiles $\mathbb{R}^{n}.$
3. NOTION OF PRE-TILES AND TOPS
Figure 3. This picture relates to the IFS in Equation 3.3 and shows successive tilings of the form $f_{-\mathbf{k}|n}(\{\overline{A_{\mathbf{i}|(n+1)}}:\mathbf{i\in\Sigma\})}$ where $\mathbf{k}|4=1414.$ In successive panels the bounding box shown in red remains constant.
Figure 4. This shows the converged tiling, within the red bounding box, following on from Figure 3. See text. As the tiling develops further, parts within the bounding box cease to change.
3.1. Basic example of top tiling when the OSC holds
Here we use a simple two-dimensional example, where the OSC holds, to illustrate notions and language associated with top tilings. We consider the IFS
(3.1) $\{\mathbb{C};f_{1}(z)=z/2,f_{2}(z)=(z+1)/2,f_{3}(z)=(z+i)/2,f_{4}(z)=(z+1+i)/2\}$
which has attractor $A=[0,1]\times[0,1].$ We may think of $A$ as being tiled by the four half-sized copies $\left\{f_{1}(A),f_{2}(A),f_{3}(A),f_{4}(A)\right\}$ of $A.$ Each point of $A$ belongs to a unique pre-tile, specifically
$A_{1}=[0,\frac{1}{2}]\times[0,\frac{1}{2}],A_{2}=(\frac{1}{2},1]\times[0,\frac{1}{2}],A_{3}=[0,\frac{1}{2}]\times(\frac{1}{2},1],A_{4}=(\frac{1}{2},1]\times(\frac{1}{2},1].$
See Figure 1. These pre-tiles, which generally are neither open nor closed, but have non-empty interiors, are obtained by taking the “top” of the set $\{f_{1}(A),f_{2}(A),f_{3}(A),f_{4}(A)\}$, after assigning lexicographical ordering to the indices $\{1,2,3,4\}$. Here the set $f_{1}(A)$ lies “on top of” parts of $f_{2}(A),f_{3}(A),$ and $f_{4}(A)$. What we call the “top tiling” of $A$ in this example comprises four pre-tiles, the portions of the sets  $f_{i}(A)$ that belong to the “top”.
We extend this top tiling to a larger region in two steps. (i)We use
$A={\textstyle\bigcup_{i,j=1}^{4}}f_{ij}(A)={\textstyle\bigcup_{i,j=1}^{4}}A_{ij},$
[…]
$A_{11}=[0,\frac{1}{4}]\times[0,\frac{1}{4}],A_{12}=(\frac{1}{4},\frac{1}{2}]\times[0,\frac{1}{4}],A_{13}=(\frac{1}{2},\frac{3}{4}]\times[0,\frac{1}{4}],A_{14}(A)=(\frac{3}{4},1]\times[0,\frac{1}{4}].$
That is, we have used the lexicographic ordering $11>12>13>14>21>\dots$ to determine to which tile to assign points which belong to more than one set in $\{f_{ij}(A)\}.$ Note that in general $A_{ij}\neq f_{ij}(A)$ because $f_{ij}(A)$ is closed but $A_{ij}$ is typically not closed. (ii)We apply $f_{1}^{-1}$ to $\left\{A_{ij}\right\}$ to obtain a partition of $0,2]\times[0,2]$ consisting of the sixteen disjoint pre-tiles illustrated in the right-hand panel Figure 2,
$f_{1}^{-1}A_{11}$ $=[0,0.5]\times(0,0.5],f_{1}^{-1}A_{12}=(0.5,1]\times[0,0.5],$
$f_{1}^{-1}A_{13}$ $=[0.5,0.5]\times(0,1],f_{1}^{-1}A_{14}=(0.5,2]\times[0,0.5],$
$\dots$
$f_{1}^{-1}A_{43}$ $=(1.5,2]\times(1,1.5],f_{1}^{-1}A_{44}=(1.5,2]\times(1.5,2].$
This tiling extends the pre-tiling $\{A_{i}\},$ in particular
$f_{1}^{-1}A_{1j}=A_{j}\text{ for }j=1,2,3,4.$
We consider a second extension of the top tiling in (i), this time yielding a tiling of $1,1]\times[-1,1],$ by applying $f_{4}^{-1}$ to $\left\{A_{ij}\right\},$ as illustrated in the left-hand panel in Figure 2. In this case we notice that
(3.2) $f_{4}^{-1}A_{4j}\subsetneq A_{j}\text{ for }j=1,2,3.$
This example may be developed further to yield obvious tilings, by squares, of the plane, half-plane and quarter plane, depending on the choice of $\mathbf{j}$, typifying aspects of self-similar polygonal tilings, as discussed in [8], provided by the formula
$\Pi(\mathbf{j})=\lim_{n\rightarrow\infty}f_{-\mathbf{j}|k}\{f_{\mathbf{i}|k+1}(A):\mathbf{i\in}\Sigma\}.$
A similar development applies to any IFS that obeys the OSC. See for example [8]. Our main purpose in this simple example is to draw attention to Equation 3.2 and the notion of pre-tiles.
Figure 5. The overlapping attractor of an IFS of two similitudes, each with the same scaling factor. The figure has been rotated ninety degrees anticlockwise relative to the usual presentation of the x,y coordinate system.
3.2. Illustrations of pre-tiles when the OSC does not hold
We consider the overlapping IFS
(3.3) $\{\mathbb{C};f_{1}(z)=\frac{3}{5}z,f_{2}(z)=\frac{3z+2}{5},f_{3}(z)=\frac{3z+2i}{5},f_{4}(z)=\frac{3z+2+2i}{5}\}.$
The left hand panel in Figure 3 illustrates the attractor, namely the unit square, with the four pre-tiles
$A_{1}$ $=[0,0.6]\times[0,0.6],A_{2}=(0.6,1]\times[0,0.6],$
$A_{3}$ $=[0,0.6]\times(0.4,1],A_{4}=(0.6,1]\times(0.6,1].$
The subsequent panels illustrate successively the partitions $f_{1}^{-1}(\left\{A_{i_{1}i_{2}}:i_{1}i_{2}\in\Sigma_{2}\right\})$, $f_{1}^{-1}f_{4}^{-1}(\left\{A_{i_{1}i_{2}i_{3}}:i_{1}i_{2}i_{3}\in\Sigma_{3}\right\},f_{1}^{-1}f_{4}^{-1}f_{1}^{-1}(\left\{A_{i_{1}i_{2}i_{3}i_{4}}:i_{1}i_{2}i_{3}i_{4}\in\Sigma_{4}\right\}).$ In each panel the same viewing window is illustrated in red. It can be seen that in successive panels, pre-tiles that cover approximately the same region either stay the same or shrink, while new pre-tiles appear. Figure 4 illustrates a close-up of the part of the partition $f_{1}^{-1}f_{4}^{-1}f_{1}^{-1}f_{4}^{-1}(\left\{A_{i_{1}i_{2}i_{3}i_{4}i_{5}}:i_{1}i_{2}i_{3}i_{4}i_{5}\in\Sigma_{5}\right\})$ that lies within the bounding box. This patch of the final tiling has now converged, and is a subset of $f_{-\mathbf{k}|n}(\left\{A_{\mathbf{i|}n+1}:\mathbf{i}\in\Sigma\right\})$ for all $\mathbf{k}$ such that $\mathbf{k}|4=1414.$
4. EXAMPLES OF TOPS TILINGS
Figure 6. Illustration of the top of the leaf attractor at depths 0 to 5.
4.1. Leaf example of a two-dimensional top tiling
Figure 7. This shows the sequence of tops $\Pi(111...|n)$ for $n=0,1,\dots,6$ for the leaf IFS. In each case the tip of the stem is at the same point in the xy-plane.
We consider the IFS defined by the two similitudes
(4.1) $f_{1}(x,y)$ $=(0.7526x-0.2190y+0.2474,0.2190x+0.7526y-0.0726)$
$f_{2}(x,y)$ $=(-0.7526x+0.2190y+1.0349,0.2190x+0.7526y+0.0678)$
The overlapping attractor looks like a leaf, and is the union of two overlapping copies of itself, as illustrated in Figure 5. The point with address $\overline{\mathbf{1}}=111\dots$ is represented by the tip of the stem of the leaf. The stem is actually arranged in an infinite spiral, not visible in the picture.
Figure 6 illustrates the top of $L$ at depths $n\in\{0,1,\dots,5\}$ labelled for the most part by the addresses in $\Sigma_{n}$ .
Figure 7 illustrates the successive top tilings $\Pi(\overline{1}|n)$ for $n=0,1,\dots,5$ for the IFS in Equation 4.1. In this case the pre-tiles do not shrink for some initial set of levels because $\overline{1}$ is the lexicographically maximal address. The pictures represent successively larger fully converged parts of the unbounded tiling $\Pi(\overline{1}).$ They were computed by the same method as those in Figure 8.
Figure 8. Example of a sequence of pre-tilings associated with a non-reversible address: it illustrates the part of $\Pi(\overline{21}|n)$ that intersects the square $0,1]\times[0,1],$ for $n=0,1,2,3,4,5,6,7.$ See text.
Figure 8 illustrates the part of $\Pi(\overline{21}|n)$ that intersects the square $0,1]\times[0,1],$ for $n=0,1,2,3,4,5,6,7.$ These illustrations were computed using the chaos game algorithm, assigning unique colours ”stolen” from a master image, to identify pre-tiles by their addresses to depth $n,$ illustrated in the bottom eight panels, then using an edge detection routine. While it appeared from these pictures that $\overline{21}\in\overleftarrow{\Sigma},$ we found later that $\Pi(\overline{21}|8)$ contains a new tiles in $0,1]\times[0,1]$ that overlap many tiles in $\Pi(\overline{21}|6)$. A subsequent calculation convinced us that $\overline{21}\notin\overleftarrow{\Sigma}.$
However Figure 9 correctly represents the part of the top tiling $\Pi(\overline{2})$ that lies within $0,1]\times[0,1].$ In this case it is readily checked that $\overline{2}\in\overleftarrow{\Sigma}.$
Figure 9. See text. This illustrates the part of the tiling $\Pi(\overline{2})$ within the window $0,1]\times[0,1].$
Figure 10. Within the same window, two of many different top tilings associated with an aperiodic monotile of Smith et al. The top tiling of the attractor of the IFS is illustrated in green. See text.
4.2. Top tilings related to hat tilings
Figure 10 illustrates the attractor and the part of each of the top tilings $\Pi(\overline{5324})$ and $\Pi(\overline{2345})$ for the following IFS, within the viewing window $1\leq x\leq 2,$ $1.5\leq y\leq 1.5.$ The tilings differ in the top left quadrant.
$f_{1}(z)=\frac{2}{\sqrt{5}+3}z$
$f_{2}(z)=\frac{2}{\sqrt{5}+3}e^{(-i\pi/3)}z+\frac{1}{4}\left(\sqrt{5}-1\right)\left(i\sqrt{3}+1\right)$
$f_{3}(z)=\frac{2}{\sqrt{5}+3}e^{(-i2\pi/3)}z+\frac{1}{4}\left(\sqrt{5}-1\right)\left(i\sqrt{3}+3\right)$
$f_{4}(z)=\frac{2}{\sqrt{5}+3}e^{(2i\pi/3)}z+\frac{1}{4}\left(i\sqrt{3}+\frac{1}{4}\sqrt{5}\right)+\frac{3}{4}-\frac{1}{4}i\sqrt{15}$
$f_{5}(z)=\frac{2}{\sqrt{5}+3}e^{(i\pi/3)}z+\frac{1}{4}(i\sqrt{3}-\sqrt{5})+\frac{5}{4}-\frac{1}{4}i\sqrt{15}$
$f_{6}(z)=\frac{2}{\sqrt{5}+3}z+\frac{1}{2}(3-\sqrt{5})$
$f_{7}(z)=\frac{2}{\sqrt{5}+3}z+\frac{1}{2}(\sqrt{5}-1)$
This IFS was derived from study of Figure 2.14 in [19]. It provides an alternative description of structure of the limiting fractal tiling shapes associated with hat tilings. Instead of there being two different structures involved there is a single shape. The tile corresponding to $f_{7}(z),$ namely $f_{7}(A),$ has a missing piece because it lies partly underneath $f_{6}(A).$ Otherwise the sets $f_{i}(A)$ are just-touching. In fact this system is of finite type and may be represented by a graph directed IFS. Here different, non-homeomorphic tilings are codified by their corresponding reversible addresses, namely the infinite labelled paths of the transposed corresponding graph. See Figure 10. The tiled attractor of the IFS is indicated in green. More details concerning this example are included in [12].
5. Generalization
The theory developed in this paper applies in the setting of point-fibred IFS, as defined in [7] based on the original work of [16]. In particular, top tilings may be developed in the context of loxodromic Mõbius transformations [22].
A version of the theory also applies when the tiles are defined using regions yielded by applying $f_{-\mathbf{j}|n}$ to all images of the boundary of $A$ in $f_{\mathbf{i}|n}(A)$ at depth $n$ for all $\mathbf{i}.$ In this setting Baire’s Theorem provides well-defined tilings of say $\mathbb{R}^{2},$ one for each of a dense set of choices for $\mathbf{j\in}\left\{1,2,\dots,M\right\}^{\mathbb{N}}.$ This allows tilings to be well-defined in the case of non-algebraic data, possessing some level of self-similarity, and maybe suitable for modelling images derived from nature.
6. Acknowledgement
We thank Louisa Barnsley for help with the artwork.